How do I solve the system using addition/elimination method and show work?
3x+2y=14
3x-2y+10

Found 2 solutions by dabanfield, richwmiller:
Answer by dabanfield(803)

How do I solve the system using addition/elimination method and show work?
1.) 3x+2y=14
2.) 3x-2y+10
We can subtract the corresponding sides of the two equations above to eliminate x as follows:
(3x+2y)-(3x-2y) = 14-10
3x-3x + 2y+2y = 4
4y = 4
y = 1
Now substitute 1 for y in either 1.) or 2.) and solve for x.

Answer by richwmiller(17219)

assuming you mean
3x-2y=10
3x+2y=14
add
6x=24
x=4
3*4-2y=10
12-2y=10
-2y=-2
y=1
(4,1)